OpenStudy (anonymous):

A car is traveling at 30 m/s when the driver slams on the brakes. From the that point, the car stops in 10 meter. How long did it take the car to stop from when the brakes were applied? Can you solve this question?

OpenStudy (ghazi):

use \[v=u-a*t\] i have used a in negative because it is retarding ... now final velocity=0 and U=30

OpenStudy (ghazi):

also s=ut-0.5*a*t^2

OpenStudy (ghazi):

10=30*t-0.5*a*t^2

OpenStudy (ghazi):

and (30)^2= 2*a*10